OpenStudy (anonymous):

cos^(2)x(1+tan^(2)x)=1 please help prove this

OpenStudy (anonymous):

(1-cosx)(1+cosx)=sin^(2)x ??

OpenStudy (zpupster):

\[\cos ^{2}x + \cos ^{2}x(\sin ^{2}x)/\cos ^{2} x\] we have multiplied the cos^2 x thru and tan ^2x = sin^2x/cos^2x cancelling the cos^2x 's out that leaves the magic identity: sin^2x + cos^2x = 1

OpenStudy (anonymous):

ohh i see !!! Thanks !! How would you do the bottom one?

OpenStudy (anonymous):

@Loser66 do you know how to do this? 1-cosx)(1+cosx)=sin^(2)x

OpenStudy (loser66):

(a+b)(a-b)= a^2-b^2 yours is (1-cos)(1+cos) = 1-cos^2 =sin^2 , quite clear, right?

OpenStudy (anonymous):

lost

OpenStudy (loser66):

why? a is 1 b is cos

OpenStudy (anonymous):

so im going to use the Pythagorean therom?

OpenStudy (loser66):